if log 0.57 =1.756 Then value if log57+log (.57)^3+log√.57
 

Grade:12th pass

2 Answers

UTTAM SAHA
5 years ago
First express  the log in terms of 0.57 as shown below
log57 + log(.57)^3+log\sqrt{.57}
=log($.57\times10$) + log(.57)^3+log(.57)^{1/2}
=log.57+log10 + 3 log(.57)+\frac{1}{2}log(.57)
Now log 0.57 =1.756 given and log 10=1 then 
 
=1.756+1+3x1.756+0.5x1.756
=8.902 This is the answer.
UTTAM SAHA
5 years ago
Hey I made a small mistake
log57=log(0.57x100)=log(0.57)+log100=log(0.57)+log10^2=log(0.57)+2 log10
please correct it. The result will be 9.902